OpenStudy (jkbo):

Please help You are given the coordinates of the vertex (8,8) and of a point (5,1) on a parabola. Find the coefficients for the equation of the parabola.

OpenStudy (anonymous):

do you know the vertex form?

OpenStudy (jkbo):

No I dont

OpenStudy (anonymous):

y = a(x-h)^2 + k, where (h,k) is the vertex and a is non-zero constant. What is (h,k) in this case?

OpenStudy (jkbo):

(8,8)

OpenStudy (anonymous):

if you plug that in the formula, what do you get?

OpenStudy (jkbo):

y=a(x-8)^2+8

OpenStudy (jkbo):

@sourwing I'm not sure what to do from here

OpenStudy (phi):

they tell you point (5,1) is on the curve. replace x with 5, and y with 1 and solve for a

OpenStudy (anonymous):

now you need to find a. Which is why another point is given (x,y) = (5,1)

OpenStudy (jkbo):

Alright thank you both but I need to Find the coefficients for the equation of the parabola I just have to answer like this y = x2 + x + . Basicly fill in the blanks

OpenStudy (phi):

First, are you able to find a using 1 = a(3-8)^2 +8 ?

OpenStudy (jkbo):

yes Im @phi

OpenStudy (jkbo):

-7/9

OpenStudy (phi):

ok, that means your equation is \[ y = -\frac{7}{9}(x-8)^2 +8 \] to put this in standard form, expand (x-8)^2 then distribute the -7/9

OpenStudy (jkbo):

@phi Im not sure on how to expand this

OpenStudy (phi):

(x-8)(x-8) do you know FOIL? First, outer , inner , last?

OpenStudy (phi):

or use the distributive law if it were A*(x-8) you would do Ax - A*8 now if A was (x-8) this means Ax - A8= (x-8)x - (x-8)*8 now distribute again.

OpenStudy (jkbo):

alright thank you

OpenStudy (tkhunny):

OR - and this is a REALLY BIG "or" \((y-8)^{2} = -\dfrac{49}{3}(x-8)\) There just isn't enough information to determine if Vertical or Horizontal orientation is intended. There are TWO solutions with Axis of Symmetry parallel to a coordinate axis.

OpenStudy (jkbo):

@tkhunny I just have to answer like this y = x2 + x + . Basicly fill in the blanks

OpenStudy (tkhunny):

You didn't SAY that. That information rules out the Horizontal Axis of Symmetry. If you don't say it, you didn't say it. Wow! Can I copyright that?

OpenStudy (phi):

if you multiply out \( (x -8)^2 \) you should get \[ (x -8)^2 =(x-8)(x-8)\\=x^2 -8x -8x +64\\= x^2-16x +64\] that means your equation \[ y = -\frac{7}{9}(x-8)^2 +8 \\y= -\frac{7}{9}\left(x^2-16x +64\right)+8 \] now distribute the -7/9 (which means multiply each term inside the parens by -7/9 and then drop the parens)